OpenStudy (anonymous):

what is the equation of the line passing through (-3,6) and parallel to the line y=4x-2 in standard form

OpenStudy (anonymous):

as the lines are parallel their slopes are equal, meaning m=4 for both the lines. now using point-slope form of a straight line eq. which is, \[y-y _{1}=m(x-x _{1})\] you get, y-6=4(x+3) => y=4x+12+6=4x+18

OpenStudy (anonymous):

did you understand?